OpenStudy (anonymous):

8 x = 3 x − 2 (a) When you multiply the left side of the equation by the LCM x(x - 2), the left side simplifies to (b) When you multiply the right side of the equation by the LCM x(x - 2), the right side simplifies to

OpenStudy (anonymous):

8/x=3/x-2

OpenStudy (anonymous):

(a) asks what the left side is when you mutiply by x(x-2). x(x-2) * 8/x = 8(x-2), since x's cancel each other out. = 8x - 16